Complete the following statement. The electron volt is a unit of Selected Answer. Electric Potential Answers: Energy Electric Potential Charge of Proton Charge of Electron
Answer #1

Electron volt is the unit of energy.

1 eV = 1.602 × 10^(-12) erg = 1.602×10^(-19) joules

This term is usually used in nuclear physics.

electron volt is energy aquired by an electron.